Two positive point charges 4micro c and 1 micro c are at a distance of 2 m.find the point on the line joining the charges where the net electric field is 0. 

Grade:12

1 Answers

Khimraj
5 years ago
let distance from 4 micro C is x
then
(K*4*10-6)/x2 = (K*1*10-6)/(2-x)2
then
4/x2 = 1/(2-x)2
taking square root both side
2/x = 1/(2-x)
4 – 2x = x
SO x = 4/3 m from 4 micro C.
Hope it clears. If you like anser then please approve it.